This is a Biochemistry 200 level question, please answer neatly and as soon as possible. Thanks in advance, and I will like as soon as answer is posted!
The enzyme happyase catalyzes the following reaction: SAD — HAPPY An enzyme kinetics experiment gave the data shown in the taQUESTION 6 What will be the reaction rate at 16.8 mM SAD and 25.00 uM happyase?
Answer #1

for an enzyme catalysis reaction Rate = K [o] Eso] - u KM+Cso]. where, Kg = constanı [E ] = Initial cone of enzyme [s ] = subKa+C50] BEJ Cse] TRTO - Bikan tesc) » tag = tegasus ! WALE - cu tu ke + Cel. in the form y=mx+c (straight une) c = tid = Inte2 024142 slope - = km = 0.0286 Intercopr = 8.44 KIE]. - Intercept = 0.0286 - 34.965 & go, from ea ů Rate = 34.965 tefte um X